# Phonon assisted light absorption and emission in cubic-Boron Nitride

### v1

@@ Acknowledgements.
C.A. and A.P. acknowledge B. Demoulin and A. Saul for the management of the computer cluster Claudia. F.P. acknowledges Muralidhar Nalabothula for insightful discussions on exciton symmetries and the evaluation of luminescence spectra. C.A. acknowledges funding from European Research Council MSCA-ITN TIMES under grant agreement 101118915. E.C. acknowledges ANR project COLIBRI No. ANR-22-CE30-0027. C.A., A.M. and A.P. acknowledge support from the Amidex foundation through the project INDIGENA. F.P. Acknowledges support from ICSC - Centro Nazionale di Ricerca in High Performance Computing, Big Data and Quantum Computing - funded by the European Union through the Italian Ministry of University and Research under PNRR M4C2I1.4 (Grant No. CN00000013). This work has also been supported by the European Union through the MaX ”MAterials design at the eXascale” Centre of Excellence (Grant agreement No. 101093374, co-funded by the EuroHPC JU).References
- Cazorla and Gould (2019) C. Cazorla and T. Gould, Science Advances 5, eaau5832 (2019), https://www.science.org/doi/pdf/10.1126/sciadv.aau5832 .
@@
- Paleari and Marini (2022) F. Paleari and A. Marini, Phys. Rev. B 106, 125403 (2022).
- Stefanucci and Perfetto (2025) G. Stefanucci and E. Perfetto, SciPost Phys. 18, 009 (2025).
@@
- Togo et al. (2024) A. Togo, K. Shinohara, and I. Tanaka, Science and Technology of Advanced Materials: Methods 4, 2384822 (2024), https://doi.org/10.1080/27660400.2024.2384822 .
- Goodlett et al. (2024) S. M. Goodlett, N. L. Kitzmiller, J. M. Turney, and I. Schaefer, Henry F., The Journal of Chemical Physics 161, 024107 (2024), https://pubs.aip.org/aip/jcp/article-pdf/doi/10.1063/5.0216738/20039263/024107_1_5.0216738.pdf .

### v2

@@ Acknowledgements.
C.A. and A.P. acknowledge B. Demoulin and A. Saul for the management of the computer cluster Claudia. F.P. acknowledges M. Nalabothula for insightful discussions on exciton symmetries and luminescence. C.A. acknowledges funding from European Research Council MSCA-ITN TIMES under grant agreement 101118915. E.C. acknowledges ANR project COLIBRI No. ANR-22-CE30-0027. C.A., A.M. and A.P. acknowledge support from the Amidex foundation through the project INDIGENA. F.P. Acknowledges support from ICSC-CNR in High Performance Computing, Big Data and Quantum Computing - funded by the European Union(EU) through the Italian Ministry of University and Research under PNRR M4C2I1.4 (Grant No. CN00000013). This work has also been supported by the EU through the MaX ”MAterials design at the eXascale” Centre of Excellence (Grant agreement No. 101093374, co-funded by the EuroHPC JU).References
- Cazorla and Gould (2019) C. Cazorla and T. Gould, Science Advances 5, eaau5832 (2019).
@@
- Nalabothula et al. (2025) M. Nalabothula, D. Sangalli, F. Paleari, S. Reichardt, and L. Wirtz, “Symmetries of excitons,” (2025), arXiv:2511.21540 [cond-mat.mes-hall] .
@@
- Togo et al. (2024) A. Togo, K. Shinohara, and I. Tanaka, Science and Technology of Advanced Materials: Methods 4, 2384822 (2024).
- Goodlett et al. (2024) S. M. Goodlett, N. L. Kitzmiller, J. M. Turney, and I. Schaefer, Henry F., The Journal of Chemical Physics 161, 024107 (2024).
@@
- Shishkin and Kresse (2007) M. Shishkin and G. Kresse, Physical Review B—Condensed Matter and Materials Physics 75, 235102 (2007).